Find the point on the plane z = 8 – 3x + 2y for which the function x2 + 3y2 + Z is a minimum.
Answer #1

Coln - 8. 3x+2y and Let function fcx,y)= x² + 3y² + 2 Non pot z from ean. I in can 2:- . feyn x² + 3y² + 8-3x+2y ③ Let we useput (x,y in Canl to get 2- 2. 8-3x+2y = 8-34 3 + 2X – - 8-9 - - 2 3 - 48-27-4 = 214 - 12 Menco point for which function 18 mi